OpenStudy (anonymous):

find the inclination of the line passing through the points (-7,-1) (-1,2)

OpenStudy (anonymous):

2-(-1) / -1-(-7) = 3 / 6 = 1/2

OpenStudy (anonymous):

yes that is the slope, then you must find the inclination using the formula m=tan\[\theta\]

OpenStudy (anonymous):

but, im just confused whether or not i need to subtract my answer from m=tan(theta) from pi

OpenStudy (anonymous):

you mean the angle the line makes with the horizontal?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

that's arctan (1/2) on your calculator it 's probably written tan-1 26.6 degrees

OpenStudy (anonymous):

ok. thankss

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

plz close the post by clicking on Good Answer